# Accelerating Quantum Reinforcement Learning with a Quantum Natural Policy Gradient Based Approach

### v1

@@ 4.1 Outer Loop Analysis
Lemma 3 (Corollary 1 of (Mondal and Aggarwal, 2024)).
@@
$\begin{split}&J_{\rho}^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[J_{\rho}(\theta_{k})]\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+\frac{G}{K}\sum_{k=0}^{K-1}\mathbb{E}\|(\mathbb{E}\left[\omega_{k}|\theta_{k}\right]-\omega^{*}_{k})\|\\ &+\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\\ &+\dfrac{G^{2}}{\mu_{F}^{2}K}\left(\dfrac{B}{1-\gamma}+4L\mathbb{E}_{s\sim d_{\rho}^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\right)\end{split}$ (25)
@@
Lemma 3 decomposes the optimality gap in a way that achieves the order-optimal sample complexity of $\mathcal{O}(\epsilon^{-2})$ in the classical setting. Notice that both the first-order approximation error, $\mathbb{E}\|(\mathbb{E}[\omega_{k}|\theta]-\omega_{k}^{*})\|$ and the second-order error, $\mathbb{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}$ are both dependent on $H$, the number of iterations in the inner loop.
@@ 4.2 Inner Loop Analysis
$||\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||\leq\hat{\delta}_{g}$ (26)
$\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\nabla_{\theta}J_{\rho}(\theta)||\leq\frac{\sqrt{d}G}{(1-\gamma)^{2}}+\hat{\delta}_{g}$ (27)
$||\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||\leq\hat{\delta}_{F}$ (28)
$\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-F_{\rho}(\theta)||\leq\sqrt{d}G^{2}+\hat{\delta}_{F}$ (29)
@@
Analyzing the error bounds $||\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\mathbb{E}[g(\tau|\theta)]||$ and $||\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]-\mathbb{E}[F(\tau|\theta)]||$ quantifies the truncation error, showing that the truncation introduces an exponentially decaying bias of order $\operatorname*{\mathcal{O}}(\gamma^{N})$ while maintaining a bounded variance. This is achieved by leveraging properties of the geometric series and analyzing the dependence on the discount factor $\gamma$. Given the above, we can formally state the bias and variance bounds for the gradient estimator $\tilde{g}_{h}$ and Fisher estimator $\tilde{F}_{h}$ for each $h$, along with expected number of quantum samples $|\psi_{F}^{\theta}\rangle$ and $|\psi_{g}^{\theta}\rangle$ used in (18) and (19).
@@
For any $\theta$ and h, (18) and (19) satisfy $\mathbb{E}[\tilde{g}_{h}]=\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]$ and $\mathbb{E}[\tilde{F}_{h}]=\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]$ with reduced variance $\emph{Var}(\tilde{g}_{h})=\hat{\sigma}^{2}_{g}$ and $\emph{Var}(\tilde{F}_{h})=\hat{\sigma}^{2}_{F}$. Moreover, (18) and (19) perform $\tilde{\mathcal{O}}\big{(}\frac{G^{2}d}{\hat{\sigma}_{F}}\big{)}$ queries of $\mathcal{U}_{F}$ and $\tilde{\mathcal{O}}\big{(}\frac{Gd}{(1-\gamma)^{2}\hat{\sigma}_{g}}\big{)}$ queries of $\mathcal{U}_{g}$ in expectations.
@@
$||\mathbb{E}[\tilde{g}_{h}]-\nabla_{\theta}J_{\rho}(\theta)||^{2}\leq\hat{\delta}_{g}^{2}$ (31)
$\mathbb{E}\big{[}||\tilde{g}_{h}-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}\leq\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}$ (32)
$||\mathbb{E}[\tilde{F}_{h}]-F_{\rho}(\theta)||^{2}\leq\hat{\delta}_{F}^{2}$ (33)
$\mathbb{E}\big{[}||\tilde{F}_{h}-F_{\rho}(\theta)||^{2}\big{]}\leq\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}$ (34)
@@
Note that $\hat{\delta}_{g}^{2}$ and $\hat{\delta}_{F}^{2}$ in Lemma 5 characterizes the exponentially decaying bias by truncation. The proof of Lemma 5 (in Appendix D) decomposes the second order error into the sum of bias and variance terms, which are futher bounded by Theorem 1 and Theorem 2 respectively. Compared with the classical NPG algorithms in (Mondal and Aggarwal, 2024; Liu et al., 2020), the extra bias terms requires a more careful analysis. We to derive the error bound for the first-order and second-order approximation of $\omega_{k}^{*}$ as follows:
@@
$\mathbb{E}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\leq\exp\left(-{H\alpha\mu_{F}}\right)\mathbb{E}\|\omega_{0}-\omega^{*}\|^{2}+C_{0}$ (35)
@@
$\|\mathbb{E}[\omega_{k}]-\omega_{k}^{*}\|^{2}\leq\exp\left(-{H\alpha\mu_{F}}\right)\|\mathbb{E}[\omega_{0}]-\omega^{*}\|^{2}+C_{1}$ (36)
@@
$\begin{split}C_{0}&=4\mu_{F}^{-2}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha\mu_{F}^{-1}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]={\operatorname*{\mathcal{O}}}\left(\gamma^{2N}+\frac{\hat{\sigma}_{F}^{2}}{(1-\gamma)^{4}}+\hat{\sigma}_{g}^{2}\right)\end{split}$ (37)
$\begin{split}C_{1}&=\dfrac{6}{\mu_{F}}\left(\alpha+\dfrac{1}{\mu_{F}}\right)\bigg{[}\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\bigg{]}={\operatorname*{\mathcal{O}}}\left(\gamma^{2N}\right)\end{split}$ (38)
@@
The proof of Lemma 6 (in Appendix E) shows that the bias terms $\operatorname*{\mathcal{O}}(\gamma^{N})$ from the gradient and Fisher matrix estimators would accumulate across iterations, leading to a total error contribution of the form
@@ 4.3 Final Result
$J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[J_{\rho}(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\epsilon$ (39)
@@
The resulting sample complexity $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ is due to the quantum speedup in the inner loop while the outer loop maintained as in the classical settings. This aligns with the results of quantum stochastic optimizations (Sidford and Zhang, 2024) due to the same double-loop algorithm structure with classical outer loop and quantum inner loop. Note that $H$ is chosen to be logarithmically dependent on the total samples to account for the exponentially decaying bias term of the truncation error.
@@ Appendix B Details of Quantum Variance Reduction
Algorithm 2 QVarianceReduce (Algorithm 2 in Sidford and Zhang (2024))
@@
2:Output: An estimate $\hat{\mu}$ satisfying $\mathbb{E}{\|\hat{\mu}-\mathbb{E}[X]\|}^{2}\leq\hat{\sigma}^{2}$
@@
5:if ${\|X_{1}-X_{2}\|}\leq D$ then
6:     return $X_{1}$
@@
Lemma 7 (Theorem 4 of (Sidford and Zhang, 2024)).
@@ Appendix C Proof of Theorem 1
where $\tau=(s_{0},a_{0},\ldots,s_{k},a_{k},\ldots)$ is an infinite length trajectory, generated by taking policy $\pi_{\theta}$ with $s_{0}$ sampled from $\rho$. (40) is a standard expression of the policy gradient unbiased estimator used in (Baxter and Bartlett, 2001; Liu et al., 2020). To prove that (41) is an unbiased estimator of $F_{\rho}(\theta)$, note that
@@
$||\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||$ $=||\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\mathbb{E}[g(\tau|\theta)]||$
$=\bigg{|}\bigg{|}\mathbb{E}\bigg{[}\sum_{n=N}^{\infty}\left(\sum_{t=0}^{n}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right)\left(\gamma^{n}r(s_{n},a_{n})\right)\bigg{]}\bigg{|}\bigg{|}$
@@
$\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\nabla_{\theta}J_{\rho}(\theta)||\leq\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]||+\mathbb{E}||\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||$ (50)
@@
$\text{Var}(\hat{g}_{\rho}(\tau_{N}|\theta))=\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}\leq\frac{dG^{2}}{(1-\gamma)^{4}}$ (51)
Since $\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]||\leq\sqrt{\mathbb{E}||\hat{g}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}}$ due to Jensen’s inequality, combining (49), (50) and (51) we obtain (27). Similarly,
$||\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||$ $=||\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]-\mathbb{E}[F(\tau|\theta)]||$
$=\bigg{|}\bigg{|}\mathbb{E}\bigg{[}(1-\gamma)\sum_{n=N}^{\infty}\left(\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\otimes\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\gamma^{n}\right)\bigg{]}\bigg{|}\bigg{|}$
@@
$\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-F_{\rho}(\theta)||\leq\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]||+\mathbb{E}||\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||$ (53)
@@
$\text{Var}(\hat{F}_{\rho}(\tau_{N}|\theta))=\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]||^{2}\leq dG^{4}$ (54)
Since $\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]||\leq\sqrt{\mathbb{E}||\hat{F}_{\rho}(\tau_{N}|\theta)-\mathbb{E}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}}$ due to Jensen’s inequality, combining (52), (53) and (54) we obtain (29).

### v3

@@ 4.1 Outer Loop Analysis
Lemma 3 (Corollary 1 of (Mondal & Aggarwal, 2024a)).
@@
$\begin{split}&J_{\rho}^{*}-\frac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J_{\rho}(\theta_{k})]\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+\frac{G}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}\left[\omega_{k}|\theta_{k}\right]-\omega^{*}_{k})\|\\ &+\dfrac{B}{4L}\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\\ &+\dfrac{G^{2}}{\mu_{F}^{2}K}\left(\dfrac{B}{1-\gamma}+4L\operatorname{\mathbb{E}}_{s\sim d_{\rho}^{\pi^{*}}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\right)\end{split}$ (25)
@@
Lemma 3 decomposes the optimality gap in a way that achieves the order-optimal sample complexity of $\mathcal{O}(\epsilon^{-2})$ in the classical setting. Notice that both the first-order approximation error, $\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}[\omega_{k}|\theta]-\omega_{k}^{*})\|$ and the second-order error, $\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}$ are both dependent on $H$, the number of iterations in the inner loop.
@@ 4.2 Inner Loop Analysis
$||\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||\leq\hat{\delta}_{g}$ (26)
$\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\nabla_{\theta}J_{\rho}(\theta)||\leq\frac{\sqrt{d}G}{(1-\gamma)^{2}}+\hat{\delta}_{g}$ (27)
$||\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||\leq\hat{\delta}_{F}$ (28)
$\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-F_{\rho}(\theta)||\leq\sqrt{d}G^{2}+\hat{\delta}_{F}$ (29)
@@
Analyzing the error bounds $||\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\operatorname{\mathbb{E}}[g(\tau|\theta)]||$ and $||\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]-\operatorname{\mathbb{E}}[F(\tau|\theta)]||$ quantifies the truncation error, showing that the truncation introduces an exponentially decaying bias of order $\operatorname*{\mathcal{O}}(\gamma^{N})$ while maintaining a bounded variance. This is achieved by leveraging properties of the geometric series and analyzing the dependence on the discount factor $\gamma$. Given the above, we can formally state the bias and variance bounds for the gradient estimator $\tilde{g}_{h}$ and Fisher estimator $\tilde{F}_{h}$ for each $h$, along with expected number of quantum samples $|\psi_{F}^{\theta}\rangle$ and $|\psi_{g}^{\theta}\rangle$ used in (18) and (19).
@@
For any $\theta$ and h, (18) and (19) satisfy $\operatorname{\mathbb{E}}[\tilde{g}_{h}]=\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]$ and $\operatorname{\mathbb{E}}[\tilde{F}_{h}]=\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]$ with reduced variance $\emph{Var}(\tilde{g}_{h})=\hat{\sigma}^{2}_{g}$ and $\emph{Var}(\tilde{F}_{h})=\hat{\sigma}^{2}_{F}$. Moreover, (18) and (19) perform $\tilde{\mathcal{O}}\big{(}\frac{G^{2}d}{\hat{\sigma}_{F}}\big{)}$ queries of $\mathcal{U}_{F}$ and $\tilde{\mathcal{O}}\big{(}\frac{Gd}{(1-\gamma)^{2}\hat{\sigma}_{g}}\big{)}$ queries of $\mathcal{U}_{g}$ in expectations.
@@
$||\operatorname{\mathbb{E}}[\tilde{g}_{h}]-\nabla_{\theta}J_{\rho}(\theta)||^{2}\leq\hat{\delta}_{g}^{2}$ (31)
$\operatorname{\mathbb{E}}\big{[}||\tilde{g}_{h}-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}\leq\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}$ (32)
$||\operatorname{\mathbb{E}}[\tilde{F}_{h}]-F_{\rho}(\theta)||^{2}\leq\hat{\delta}_{F}^{2}$ (33)
$\operatorname{\mathbb{E}}\big{[}||\tilde{F}_{h}-F_{\rho}(\theta)||^{2}\big{]}\leq\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}$ (34)
@@
Note that $\hat{\delta}_{g}^{2}$ and $\hat{\delta}_{F}^{2}$ in Lemma 5 characterizes the exponentially decaying bias by truncation. The proof of Lemma 5 (in Appendix D) decomposes the second order error into the sum of bias and variance terms, which are futher bounded by Theorem 1 and Theorem 2 respectively. Compared with the classical NPG algorithms in (Mondal & Aggarwal, 2024a; Liu et al., 2020), the extra bias terms requires a more careful analysis. We derive the error bound for the first-order and second-order approximation of $\omega_{k}^{*}$ as follows:
@@
$\mathbb{E}\left[\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]\leq\exp\left(-{H\alpha\mu_{F}}\right)\operatorname{\mathbb{E}}\|\omega_{0}-\omega^{*}\|^{2}+C_{0}$ (35)
@@
$\|\mathbb{E}[\omega_{k}]-\omega_{k}^{*}\|^{2}\leq\exp\left(-{H\alpha\mu_{F}}\right)\|\operatorname{\mathbb{E}}[\omega_{0}]-\omega^{*}\|^{2}+C_{1}$ (36)
@@
$\begin{split}C_{0}&=4\mu_{F}^{-2}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]\\ &\quad+6\alpha\mu_{F}^{-1}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\\ &={\operatorname*{\mathcal{O}}}\left(\gamma^{2N}+\frac{\hat{\sigma}_{F}^{2}}{(1-\gamma)^{4}}+\hat{\sigma}_{g}^{2}\right)\end{split}$ (37)
$\begin{split}C_{1}&=\dfrac{6}{\mu_{F}}\left(\alpha+\dfrac{1}{\mu_{F}}\right)\bigg{[}\hat{\delta}_{F}^{2}\bigg{\{}\mathbb{E}\left[\|\omega_{0}-\omega_{k}^{*}\|^{2}\right]\\ &\quad+\alpha R_{0}+R_{1}\bigg{\}}+\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\hat{\delta}_{F}^{2}+\hat{\delta}_{g}^{2}\bigg{]}\\ &={\operatorname*{\mathcal{O}}}\left(\gamma^{2N}\right)\end{split}$ (38)
@@
The proof of Lemma 6 (in Appendix E) proceeds by first expressing the error between the estimated gradient and the ideal gradient as a recursive update from the inner-loop iteration. It then demonstrates that each update contracts the error exponentially, meaning the error shrinks by a fixed multiplicative factor with each iteration. Subsequently, the proof carefully bounds the additional errors introduced by truncation bias and estimator variance, leading to explicit residual terms. Finally, these contraction and residual terms are combined through a recursive argument to establish the overall error guarantees.
Lemma 6 shows that the bias terms $\operatorname*{\mathcal{O}}(\gamma^{N})$ from the gradient and Fisher matrix estimators would accumulate across iterations, leading to a total error contribution of the form
@@ 4.3 Final Result
$J_{\rho}^{*}-\dfrac{1}{K}\sum_{k=0}^{K-1}\operatorname{\mathbb{E}}[J_{\rho}(\theta_{k})]\leq\sqrt{\epsilon_{\mathrm{bias}}}+\epsilon$ (39)
@@
The resulting sample complexity $\tilde{\mathcal{O}}(\epsilon^{-1.5})$ is due to the quantum speedup in the inner loop while the outer loop maintained as in the classical settings. This aligns with the results of quantum stochastic optimizations (Sidford & Zhang, 2024) due to the same double-loop algorithm structure with classical outer loop and quantum inner loop. Note that $H$ is chosen to be logarithmically dependent on the total samples to account for the exponentially decaying bias term of the truncation error.
@@ 5 Conclusion
While our study focuses on policy gradient methods, quantum actor–critic algorithms remain largely unexplored. Recent classical results offer order-optimal sample complexity guarantees (Gaur et al., 2024; Ganesh et al., 2025), making the search for quantum speedups in this setting a key open direction. Another open direction is extending the constrained setup studied in (Mondal & Aggarwal, 2024b) to the quantum setting.
Acknowledgment
This work is supported in part by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR, and the U.S. Government assumes no liability for the contents or use thereof.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Details of Quantum Variance Reduction
Algorithm 2 QVarianceReduce (Algorithm 2 in (Sidford & Zhang, 2024))
@@
2:  Output: An estimate $\hat{\mu}$ satisfying $\operatorname{\mathbb{E}}\left\lVert\hat{\mu}-\operatorname{\mathbb{E}}[X]\right\rVert^{2}\leq\hat{\sigma}^{2}$
@@
5:  if $\left\lVert X_{1}-X_{2}\right\rVert\leq D$ then
6:     Return: $X_{1}$
@@
Lemma 7 (Theorem 4 of (Sidford & Zhang, 2024)).
@@ Appendix C Proof of Theorem 1
where $\tau=(s_{0},a_{0},\ldots,s_{k},a_{k},\ldots)$ is an infinite length trajectory, generated by taking policy $\pi_{\theta}$ with $s_{0}$ sampled from $\rho$. (40) is a standard expression of the policy gradient unbiased estimator used in (Baxter & Bartlett, 2001; Liu et al., 2020). To prove that (41) is an unbiased estimator of $F_{\rho}(\theta)$, note that
@@
$||\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||$ $=||\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\operatorname{\mathbb{E}}[g(\tau|\theta)]||$
$=\bigg{|}\bigg{|}\operatorname{\mathbb{E}}\bigg{[}\sum_{n=N}^{\infty}\left(\sum_{t=0}^{n}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right)\left(\gamma^{n}r(s_{n},a_{n})\right)\bigg{]}\bigg{|}\bigg{|}$
@@
$\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\nabla_{\theta}J_{\rho}(\theta)||\leq\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]||+\operatorname{\mathbb{E}}||\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]-\nabla_{\theta}J_{\rho}(\theta)||$ (50)
@@
$\text{Var}(\hat{g}_{\rho}(\tau_{N}|\theta))=\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}\leq\frac{dG^{2}}{(1-\gamma)^{4}}$ (51)
Since $\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]||\leq\sqrt{\operatorname{\mathbb{E}}||\hat{g}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}}$ due to Jensen’s inequality, combining (49), (50) and (51) we obtain (27). Similarly,
$||\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||$ $=||\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]-\operatorname{\mathbb{E}}[F(\tau|\theta)]||$
$=\bigg{|}\bigg{|}\operatorname{\mathbb{E}}\bigg{[}(1-\gamma)\sum_{n=N}^{\infty}\left(\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\otimes\nabla_{\theta}\log\pi_{\theta}(a_{n}|s_{n})\gamma^{n}\right)\bigg{]}\bigg{|}\bigg{|}$
@@
$\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-F_{\rho}(\theta)||\leq\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]||+\operatorname{\mathbb{E}}||\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]-F_{\rho}(\theta)||$ (53)
@@
$\text{Var}(\hat{F}_{\rho}(\tau_{N}|\theta))=\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]||^{2}\leq dG^{4}$ (54)
Since $\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]||\leq\sqrt{\operatorname{\mathbb{E}}||\hat{F}_{\rho}(\tau_{N}|\theta)-\operatorname{\mathbb{E}}[\hat{g}_{\rho}(\tau_{N}|\theta)]||^{2}}$ due to Jensen’s inequality, combining (52), (53) and (54) we obtain (29).